# Magic-wavelength nanofiber-based two-color dipole trap with sub-$\lambda/2$ spacing

### v2

@@ Abstract
In summary, we have experimentally demonstrated magic-wavelength trapping of laser-cooled atoms in a novel type of nanofiber-based two-color dipole trap. By using a blue-detuned partial standing wave and two red-detuned running waves, we create two diametral one-dimensional arrays of trapping sites for cesium atoms. In this trapping potential, the spacing of two adjacent sites is $d\simeq 0.35\lambda$, i.e. only 35 % of the resonant free-space wavelength of the cesium D2 line. We determine the lifetime of the atoms in this new trapping configuration to be $\tau=8.5\,\mathrm{ms}\pm 0.4\,\mathrm{ms}$ and perform a systematic characterization of the radial and axial trap frequencies which agree well with the expectations. This realization and characterization of a sub-$\lambda/2$ spaced standing wave nanofiber-guided dipole trap is an important step towards the experimental study of novel collective radiative effects, such as selective radiance in 1D atomic arrays.
@@ Acknowledgment
M. C. and M. S. acknowledge support from the European Commission (Marie Skłodowska-Curie IF Grants Nos. 101029304 and IF Grant No. 896957). We acknowledge funding by the Alexander von Humboldt Foundation in the framework of the Alexander von Humboldt Professorship endowed by the Federal Ministry of Education and Research, funding by the European Commission under the project DAALI (No. 899275), and by the Einstein Foundation (Einstein Research Unit on Quantum Devices). We thank Christoph Clausen, Christian Liedl, and Sebastian Pucher for developing and improving the trap potential calculations.

### v3

@@ Abstract
In summary, we have experimentally demonstrated magic-wavelength trapping of laser-cooled atoms in a novel type of nanofiber-based two-color dipole trap. By using a blue-detuned partial standing wave and two red-detuned running waves, we create two diametral one-dimensional arrays of trapping sites for cesium atoms. In this trapping potential, the spacing of two adjacent sites is $d\simeq 0.35\lambda$, i.e. only 35 % of the resonant free-space wavelength of the cesium D2 line. We determine the lifetime of the atoms in this new trapping configuration to be $\tau=8.5\,\mathrm{ms}\pm 0.4\,\mathrm{ms}$ and perform a systematic characterization of the radial and axial trap frequencies which agree well with the expectations.
This realization and characterization of a sub-$\lambda/2$-spaced standing wave nanofiber-guided dipole trap is an important step towards the experimental observation of novel collective radiative effects, such as selective radiance in 1D atomic arrays. In a next step, the filling factor of the trapping site has to be increased, from currently about $20\,\%$ to ideally close to unity. This can be achieved, by increasing the loading rate using, e.g., more efficient cooling mechanisms, such as $\Lambda$-enhanced gray molasses cooling Hsiao et al. (2018). Alternatively, even with a lower filling factor, signatures of selective radiance are predicted and should for example, show up as a significantly increased reflection of light from the ensemble. In the standard two-color nanofiber-based trap, the power reflection coefficient of a guided light field is on the order of $\beta^{2}\sim 10^{-4}$. However, already a small, defect-free array of about $10$ adjacent atoms would give rise to a reflection coefficient of close to 10 %, corresponding to an increase by three orders of magnitude Hümmer and Chang . Such arrays are highly likely to be randomly present as part of large nanofiber-trapped ensembles with moderate filling factor. By removing all atoms except those in the highest density region from the trap, signatures of selective radiance could be observed.
@@ Acknowledgment
M. C. and M. S. acknowledge support from the European Commission (Marie Skłodowska-Curie IF Grants Nos. 101029304 and IF Grant No. 896957). H. L. acknowledges funding by the European Commission (ERC Grant No. 101071882). We acknowledge funding by the Alexander von Humboldt Foundation in the framework of the Alexander von Humboldt Professorship endowed by the Federal Ministry of Education and Research, funding by the European Commission under the projects DAALI (No. 899275), and SuperWave (ERC Grant No. 101071882), as well as funding by the Einstein Foundation (Einstein Research Unit on Quantum Devices). We thank Darrick Chang and Daniel Hümmer for insightful discussion and calculations of selective radiance in our nanofiber-platform. We thank Christoph Clausen, Christian Liedl, and Sebastian Pucher for developing and improving the trap potential calculations.